# Quantum Algorithms for Matrix Operations Based on Unitary Transformations and Ancillary State Measurements

### v1

@@ 3.2 Matrix Transposition
Then, introduce a pure state $|0\rangle_{D}$ for an $m$-qubit register $D$, and an auxiliary qubit $B$ in state $|0\rangle_{B}$. The initial state is give by
$|\Phi_{0}\rangle$ $=$ $|0\rangle_{D}\otimes|\Psi\rangle\otimes|0\rangle_{B}$
$=$ $\sum_{i=0}^{N-1}\sum_{k=0}^{M-1}a_{ik}|0\rangle_{D}|i\rangle_{R}|k\rangle_{C}|0\rangle_{B}.$
@@
We use a SWAP gate to exchange the states of the $D$ and $C$ registers. The operator $W_{DC}=SWAP_{DC}\otimes|0\rangle_{B}\;{{}_{B}}\langle 0|+I_{DC}\otimes|1\rangle_{B}\;{{}_{B}}\langle 1|$ is applied to $|\Phi_{0}\rangle$, yielding
$|\Phi_{1}\rangle$ $=$ $W_{DC}|\Phi_{0}\rangle$
$=$ $\sum_{i=0}^{N-1}\sum_{k=0}^{M-1}a_{ik}|i\rangle_{D}|k\rangle_{R}|0\rangle_{C}|0\rangle_{B}.$
Step 3: Measure the auxiliary qubit $B$ in the state $|0\rangle_{B}$.
By measuring the auxiliary qubit $B$ in the state $|0\rangle_{B}$, we extract the information of the transposed matrix $S$. This gives the final state:
$|\Phi_{2}\rangle$ $=$ $|\Psi_{out}\rangle|0\rangle_{C},$ (26)
@@
$|\Psi_{out}\rangle$ $=$ $\Big{(}\sum_{i=0}^{N-1}\sum_{k=0}^{M-1}a_{ik}|i\rangle_{D}|k\rangle_{R}\Big{)}.$ (27)
Clearly, the success probability of this measurement is 1.
By the analysis above, we conclude that the overall depth of the algorithm is $O(1)$, which determined by the operator $W_{DC}$.
@@ 4 Conclusion
In all these algorithms, the final result is stored in the probability amplitude of certain quantum states. Our success probability for adding one row of a matrix to another is $\frac{G^{2}}{8}$. The success probability for swapping two rows of a matrix is $\frac{1}{24}$, while the success probability of matrix transposition is 1. The success probability for the trace calculation is $\frac{|\langle\sum_{i=0}^{N-1}a_{ii}\rangle|^{2}}{2^{3n}}$. The success probabilities either depend solely on the elements of the matrix itself, or are fixed or related to the dimensions of the matrix. The larger the dimension, the smaller the success probability. To address this challenge, amplitude amplification techniques may be employed. It is worth noting that the computational depth of these algorithms is related to the dimension of the matrices. It increases logarithmically as the matrix dimension grows. Specifically, the computational depth is $O(n)$ for elementary row operations and trace calculation, and $O(1)$ for matrix transposition.
@@ Acknowledgments
This work is being supported by the following: the National Natural Science Foundation of China (NSFC) under Grants 11761073, 12075159, and 12171044; the Academician Innovation Platform of Hainan Province. We are deeply grateful to Prof. Junde Wu for his active participation and invaluable contributions to this research.

### v2

@@ 3.2 Matrix Transpose
Then, introduce a pure state $|0\rangle_{D}$ for an $m$-qubit register $D$. The initial state is give by
$|\Phi_{0}\rangle$ $=$ $|0\rangle_{D}\otimes|\Psi\rangle$
$=$ $\sum_{i=0}^{N-1}\sum_{k=0}^{M-1}a_{ik}|0\rangle_{D}|i\rangle_{R}|k\rangle_{C}.$
@@
We use a SWAP gate to exchange the states of the $D$ and $C$ registers. The operator $SWAP_{DC}$ is applied to $|\Phi_{0}\rangle$, yielding
$|\Phi_{1}\rangle$ $=$ $SWAP_{DC}|\Phi_{0}\rangle$
$=$ $|\Psi_{out}\rangle|0\rangle_{C},$
@@
$|\Psi_{out}\rangle$ $=$ $\Big{(}\sum_{i=0}^{N-1}\sum_{k=0}^{M-1}a_{ik}|k\rangle_{D}|i\rangle_{R}\Big{)}.$ (25)
Now, the new rows and new columns of the current target matrix are enumerated by the states $D$ and $R$ respectively. Clearly, the success probability of this algorithm is 1.
Since the exchange operator $SWAP_{DC}$ only involves exchange operations when acting on different qubit pairs, its computational complexity is $O(m)$.
Remark. Formally the non-square matrix can be made a square matrix by including the proper number of zero rows or columns. In this case transposition of non-square matrix can be done as a transposition of a square matrix, i.e. by the swap operator $SWAP_{RC}$. The complexity of this operation is $O(n)$.
@@ 4 Conclusion
In all these algorithms, the final result is stored in the probability amplitude of certain quantum states. Our success probability for adding one row of a matrix to another is $\frac{G^{2}}{8}$. The success probability for swapping two rows of a matrix is $\frac{1}{24}$, while the success probability of matrix transpose is 1. The success probability for the trace calculation is $\frac{|\sum_{i=0}^{N-1}a_{ii}|^{2}}{2^{3n}}$. The success probabilities either depend solely on the elements of the matrix itself, or are fixed or related to the dimensions of the matrix. The larger the dimension, the smaller the success probability. To address this challenge, amplitude amplification techniques may be employed. It is worth noting that the computational complexity of these algorithms is related to the dimension of the matrices. It increases logarithmically as the matrix dimension grows. Specifically, the computational complexity is $O(n)$ for elementary row operations and trace calculation, and $O(m)$ for matrix transpose.
@@ Acknowledgments
This work is being supported by the following: the National Natural Science Foundation of China (NSFC) under Grants 11761073, 12075159, and 12171044; the Academician Innovation Platform of Hainan Province. We are deeply grateful to Prof. Junde Wu and Alexander I. Zenchuk for their active participation and invaluable contributions to this research.
Appendix 1. Swapping two rows of a matrix
For the quantum algorithm that swaps two rows of a matrix, we will provide a detailed explanation of the quantum algorithm that swaps the 4th and 2th rows of the following $4\times 4$ matrix $A$,
$A=\begin{pmatrix}\frac{1}{4}&&\frac{1}{16}&&\frac{3}{16}&&\frac{3}{16}\\ &&&&&&\\ 0&&\frac{1}{2}&&\frac{1}{8}&&\frac{1}{8}\\ &&&&&&\\ \frac{7}{16}&&0&&\frac{1}{4}&&0\\ &&&&&&\\ \frac{3}{16}&&\frac{3}{16}&&\frac{1}{8}&&\frac{1}{2}\end{pmatrix}$
Step 1: Construct the initial state of the entire system.
First, introduce two $2$-qubit register $R_{1}$ and $C_{1}$, which will enumerate the rows and columns of the matrix $A$, respectively. The elements of the matrix $A$ are then encoded into the following pure state:
$|\Psi_{1}\rangle$ $=$ $\frac{1}{4}|00\rangle_{R_{1}}|00\rangle_{C_{1}}+\frac{1}{16}|00\rangle_{R_{1}}|01\rangle_{C_{1}}+\frac{3}{16}|00\rangle_{R_{1}}|10\rangle_{C_{1}}$
$+\frac{3}{16}|00\rangle_{R_{1}}|11\rangle_{C_{1}}+0|01\rangle_{R_{1}}|00\rangle_{C_{1}}+\frac{1}{2}|01\rangle_{R_{1}}|01\rangle_{C_{1}}$
$+\frac{1}{8}|01\rangle_{R_{1}}|10\rangle_{C_{1}}+\frac{1}{8}|01\rangle_{R_{1}}|11\rangle_{C_{1}}+\frac{7}{16}|10\rangle_{R_{1}}|00\rangle_{C_{1}}$
$+0|10\rangle_{R_{1}}|01\rangle_{C_{1}}+\frac{1}{4}|10\rangle_{R_{1}}|10\rangle_{C_{1}}+0|10\rangle_{R_{1}}|11\rangle_{C_{1}}$
$+\frac{3}{16}|11\rangle_{R_{1}}|00\rangle_{C_{1}}+\frac{3}{16}|11\rangle_{R_{1}}|01\rangle_{C_{1}}+\frac{1}{8}|11\rangle_{R_{1}}|10\rangle_{C_{1}}$
$+\frac{1}{2}|11\rangle_{R_{1}}|11\rangle_{C_{1}}.$
Next, introduce an $2\times 2$ auxiliary matrix $Z$,
$Z=\begin{pmatrix}0&&0&&0&&0\\ &&&&&&\\ 0&&\frac{1}{\sqrt{3}}&&0&&\frac{1}{\sqrt{3}}\\ &&&&&&\\ 0&&0&&0&&0\\ &&&&&&\\ 0&&0&&0&&\frac{1}{\sqrt{3}}\end{pmatrix}$,
along with two $2$-qubit registers $R_{2}$ and $C_{2}$. The elements of $Z$ are then encoded into the following pure state:
$|\Psi_{2}\rangle=\frac{1}{\sqrt{3}}(|01\rangle_{R_{2}}|11\rangle_{C_{2}}+|11\rangle_{R_{2}}|11\rangle_{C_{2}}+|01\rangle_{R_{2}}|01\rangle_{C_{2}}).$
The two pure states $|\Psi_{1}\rangle$ and $|\Psi_{2}\rangle$ can be tensor-producted to construct the initial state of the entire system
$|\Phi_{0}\rangle$ $=$ $|\Psi_{1}\rangle\otimes|\Psi_{2}\rangle$
$=$ $\frac{1}{\sqrt{3}}(|\Psi_{1}\rangle|01\rangle_{R_{2}}|11\rangle_{C_{2}}+|\Psi_{1}\rangle|11\rangle_{R_{2}}|11\rangle_{C_{2}}$
$+|\Psi_{1}\rangle|01\rangle_{R_{2}}|01\rangle_{C_{2}})$
Step 2: Label the distinct states of $R_{2}$ and $C_{2}$.
[…]
$=$ $\frac{1}{\sqrt{3}}(|\Psi_{1}\rangle|01\rangle_{R_{2}}|11\rangle_{C_{2}}|1\rangle_{B_{1}}$
$+|\Psi_{1}\rangle|01\rangle_{R_{2}}|01\rangle_{C_{2}}|0\rangle_{B_{1}}+|\Psi_{1}\rangle|11\rangle_{R_{2}}|11\rangle_{C_{2}}|0\rangle_{B_{1}}).$
That is, label the terms where the states of $R_{2}$ and $C_{2}$ are different with $|1\rangle_{B_{1}}$, and label the terms where the states of $R_{2}$ and $C_{2}$ are the same with $|0\rangle_{B_{1}}$.
Introducing a $1$-qubit ancilla $B_{2}$ with a state of $|0\rangle_{B_{2}}$, construct the following control operator
[…]
$|\Phi_{2}\rangle=W^{(2)}_{R_{1}R_{2}C_{2}B_{2}}|\Phi_{1}\rangle|0\rangle_{B_{2}}$ (28)
$=\frac{1}{\sqrt{3}}\Big{[}\Big{(}\frac{1}{4}|00\rangle_{R_{1}}|00\rangle_{C_{1}}|01\rangle_{R_{2}}|11\rangle_{C_{2}}$
$+\frac{1}{16}|00\rangle_{R_{1}}|01\rangle_{C_{1}}|01\rangle_{R_{2}}|11\rangle_{C_{2}}$
$+\frac{3}{16}|00\rangle_{R_{1}}|10\rangle_{C_{1}}|01\rangle_{R_{2}}|11\rangle_{C_{2}}$
$+\frac{3}{16}|00\rangle_{R_{1}}|11\rangle_{C_{1}}|01\rangle_{R_{2}}|11\rangle_{C_{2}}$
$+\frac{7}{16}|10\rangle_{R_{1}}|00\rangle_{C_{1}}|01\rangle_{R_{2}}|11\rangle_{C_{2}}$
$+0|10\rangle_{R_{1}}|01\rangle_{C_{1}}|01\rangle_{R_{2}}|11\rangle_{C_{2}}$
$+\frac{1}{4}|10\rangle_{R_{1}}|10\rangle_{C_{1}}|01\rangle_{R_{2}}|11\rangle_{C_{2}}$
$+0|10\rangle_{R_{1}}|11\rangle_{C_{1}}|01\rangle_{R_{2}}|11\rangle_{C_{2}}\Big{)}|1\rangle_{B_{1}}|00\rangle_{B_{2}}$
$+\Big{(}\frac{3}{16}|11\rangle_{R_{1}}|00\rangle_{C_{1}}|01\rangle_{R_{2}}|01\rangle_{C_{2}}$
$+\frac{3}{16}|11\rangle_{R_{1}}|01\rangle_{C_{1}}|01\rangle_{R_{2}}|01\rangle_{C_{2}}$
$+\frac{2}{16}|11\rangle_{R_{1}}|10\rangle_{C_{1}}|01\rangle_{R_{2}}|01\rangle_{C_{2}}$
$+\frac{2}{4}|11\rangle_{R_{1}}|11\rangle_{C_{1}}|01\rangle_{R_{2}}|01\rangle_{C_{2}}\Big{)}|0\rangle_{B_{1}}|10\rangle_{B_{2}}$
$+\Big{(}0|01\rangle_{R_{1}}|00\rangle_{C_{1}}|11\rangle_{R_{2}}|11\rangle_{C_{2}}$
$+\frac{2}{4}|01\rangle_{R_{1}}|01\rangle_{C_{1}}|11\rangle_{R_{2}}|11\rangle_{C_{2}}$
$+\frac{2}{16}|01\rangle_{R_{1}}|10\rangle_{C_{1}}|11\rangle_{R_{2}}|11\rangle_{C_{2}}$
$+\frac{2}{16}|01\rangle_{R_{1}}|11\rangle_{C_{1}}|11\rangle_{R_{2}}|11\rangle_{C_{2}}\Big{)}|0\rangle_{B_{1}}|01\rangle_{B_{2}}\Big{]}$
$+|g_{1}\rangle_{R_{1}C_{1}R_{2}C_{2}B_{1}B_{2}}$
Step 4: Use C-SWAP gates to swap the states of $R_{1}$ and $R_{2}$, as well as the states of $R_{1}$ and $C_{2}$.
To achieve the goal of swapping the $4$ th and $2$ th rows of the matrix, apply C-SWAP gate operations to the terms in $|\Phi_{2}\rangle$ labeled by $|10\rangle_{B_{2}}$ and $|01\rangle_{B_{2}}$, swapping the states of $R_{1}$ and $R_{2}$, as well as the states of $R_{1}$ and $C_{2}$. For this, construct the following control operators:
$W^{(1)}$ $=$ $SWAP_{R_{1},C_{2}}\otimes|1\rangle_{{B_{2}}^{1}}\;{{}_{{B_{2}}^{1}}}\langle 1|$
$+I_{R_{1}C_{2}}\otimes|0\rangle_{{B_{2}}^{1}}\;{{}_{{B_{2}}^{1}}}\langle 0|$
$W^{(2)}$ $=$ $SWAP_{R_{1},R_{2}}\otimes|1\rangle_{{B_{2}}^{2}}\;{{}_{{B_{2}}^{2}}}\langle 1|$
$+I_{R_{1}R_{2}}\otimes|0\rangle_{{B_{2}}^{2}}\;{{}_{{B_{2}}^{2}}}\langle 0|$
Apply $W^{(3)}_{R_{1}R_{2}C_{2}B_{2}}=W^{(2)}W^{(1)}$ to $|\Phi_{2}\rangle$
$|\Phi_{3}\rangle$ $=$ $W^{(3)}_{R_{1}R_{2}C_{2}B_{2}}|\Phi_{2}\rangle$
$=$ $\frac{1}{\sqrt{3}}\Big{[}\Big{(}|\psi_{1}\rangle_{R_{1}C_{1}}+|\psi_{3}\rangle_{R_{1}C_{1}}\Big{)}$
$|01\rangle_{R_{2}}|11\rangle_{C_{2}}|1\rangle_{B_{1}}|00\rangle_{B_{2}}$
$+|\psi_{2}\rangle_{R_{1}C_{1}}|01\rangle_{R_{2}}|11\rangle_{C_{2}}|0\rangle_{B_{1}}|10\rangle_{B_{2}}$
$+|\psi_{4}\rangle_{R_{1}C_{1}}|01\rangle_{R_{2}}|11\rangle_{C_{2}}|0\rangle_{B_{1}}|01\rangle_{B_{2}}\Big{]}$
$+|g_{2}\rangle_{R_{1}C_{1}R_{2}C_{2}B_{1}B_{2}}$
where $|\psi_{i}\rangle_{R_{1}C_{1}},\ i=1,2,3,4$ represents the first, second, third, fourth row of the target matrix, respectively:
$|\psi_{1}\rangle_{R_{1}C_{1}}$ $=$ $\frac{1}{4}|00\rangle_{R_{1}}|00\rangle_{C_{1}}+\frac{1}{16}|00\rangle_{R_{1}}|01\rangle_{C_{1}}$
$+\frac{3}{16}|00\rangle_{R_{1}}|10\rangle_{C_{1}}+\frac{3}{16}|00\rangle_{R_{1}}|11\rangle_{C_{1}},$
$|\psi_{2}\rangle_{R_{1}C_{1}}$ $=$ $\frac{3}{16}|01\rangle_{R_{1}}|00\rangle_{C_{1}}+\frac{3}{16}|01\rangle_{R_{1}}|01\rangle_{C_{1}}$
$+\frac{2}{16}|01\rangle_{R_{1}}|10\rangle_{C_{1}}+\frac{2}{4}|01\rangle_{R_{1}}|11\rangle_{C_{1}},$
$|\psi_{3}\rangle_{R_{1}C_{1}}$ $=$ $\frac{7}{16}|10\rangle_{R_{1}}|00\rangle_{C_{1}}+0|10\rangle_{R_{1}}|01\rangle_{C_{1}}$
$+\frac{1}{4}|10\rangle_{R_{1}}|10\rangle_{C_{1}}+0|10\rangle_{R_{1}}|11\rangle_{C_{1}},$
$|\psi_{4}\rangle_{R_{1}C_{1}}$ $=$ $0|11\rangle_{R_{1}}|00\rangle_{C_{1}}+\frac{2}{4}|11\rangle_{R_{1}}|01\rangle_{C_{1}}$
$+\frac{2}{16}|11\rangle_{R_{1}}|10\rangle_{C_{1}}+\frac{2}{16}|11\rangle_{R_{1}}|11\rangle_{C_{1}}$
Step 5: Label the useful information terms and useless information terms
Introduce a 1-qubit auxiliary $B_{3}$ in the state $|0\rangle_{B_{3}}$, and use $B_{3}$ to label the useful and useless information terms in $|\Phi_{3}\rangle$. To do this, construct a control operator:
$W^{(4)}_{B_{1}B_{2}B_{3}}=V^{(1)}V^{(2)}V^{(3)},$
where $V^{(i)}=P^{(i)}_{B_{1}B_{2}}\otimes\sigma^{(x)}_{B_{3}}+(I_{B_{1}B_{2}}-P_{B_{1}B_{2}})\otimes I_{B_{3}}$, $P^{(i)}_{B_{1}B_{2}}(i=1,2,3)$ is the projection operator
$P^{(1)}_{B_{1}B_{2}}$ $=$ $|100\rangle_{B_{1}B_{2}}\;{{}_{B_{1}B_{2}}}\langle 100|$
$P^{(2)}_{B_{1}B_{2}}$ $=$ $|001\rangle_{B_{1}B_{2}}\;{{}_{B_{1}B_{2}}}\langle 001|$
$P^{(3)}_{B_{1}B_{2}}$ $=$ $|010\rangle_{B_{1}B_{2}}\;{{}_{B_{1}B_{2}}}\langle 010|.$
Apply $W^{(4)}_{B_{1}B_{2}B_{3}}$ to $|\Phi_{3}\rangle|0\rangle_{B_{3}}$, obtained:
$|\Phi_{4}\rangle$ $=$ $W^{(4)}_{B_{1}B_{2}B_{3}}|\Phi_{3}\rangle|0\rangle_{B_{3}}$
$=$ $\frac{1}{\sqrt{3}}\Big{[}\Big{(}|\psi_{1}\rangle_{R_{1}C_{1}}+|\psi_{3}\rangle_{R_{1}C_{1}}\Big{)}$
$|01\rangle_{R_{2}}|11\rangle_{C_{2}}|1\rangle_{B_{1}}|00\rangle_{B_{2}}$
$+|\psi_{2}\rangle_{R_{1}C_{1}}|01\rangle_{R_{2}}|11\rangle_{C_{2}}|0\rangle_{B_{1}}|10\rangle_{B_{2}}$
$+|\psi_{4}\rangle_{R_{1}C_{1}}|01\rangle_{R_{2}}|11\rangle_{C_{2}}|0\rangle_{B_{1}}|01\rangle_{B_{2}}\Big{]}|1\rangle_{B_{3}}$
$+|g_{2}\rangle_{R_{1}C_{1}R_{2}C_{2}B_{1}B_{2}}$
Mark the useful information terms with $|1\rangle_{B_{3}}$ and the useless information terms with $|0\rangle_{B_{3}}$.
Step 6: Use Hadamard gates for information superposition.
Apply the Hadamard operator $W^{(5)}_{B_{1}B_{2}}=H^{\otimes 3}$ acts on $B_{1}$ and $B_{2}$, obtained:
$|\Phi_{5}\rangle$ $=$ $W^{(5)}_{B_{1}B_{2}}|\Phi_{4}\rangle$
$=$ $\frac{1}{(\sqrt{2})^{3}\cdot\sqrt{3}}\Big{[}|\psi_{1}\rangle_{R_{1}C_{1}}+|\psi_{3}\rangle_{R_{1}C_{1}}+|\psi_{2}\rangle_{R_{1}C_{1}}$
$+|\psi_{4}\rangle_{R_{1}C_{1}}\Big{]}|01\rangle_{R_{2}}|11\rangle_{C_{2}}|0\rangle_{B_{1}}|00\rangle_{B_{2}}|1\rangle_{B_{3}}$
$+|g_{3}\rangle_{R_{1}C_{1}R_{2}C_{2}B_{1}B_{2}B_{3}}.$
Step 7: Measure the auxiliary $B_{1},B_{2},B_{3}$ in the state $|0001\rangle_{B_{1}B_{2}B_{3}}$.
Finally, use $|0001\rangle_{B_{1}B_{2}B_{3}}$ to measure ancilla $B_{1},B_{2},B_{3}$, thus removing useless information items and obtaining useful information items as
$|\Phi_{6}\rangle$ $=$ $|\Psi_{out}\rangle\ |01\rangle_{R_{2}}|11\rangle_{C_{2}},$
where
$|\Psi_{out}\rangle$ $=$ $|\psi_{1}\rangle_{R_{1}C_{1}}+|\psi_{2}\rangle_{R_{1}C_{1}}+|\psi_{3}\rangle_{R_{1}C_{1}}+|\psi_{4}\rangle_{R_{1}C_{1}}$
The probability of the above measurement is $\frac{1}{24}$.